Jared is 7 years older than Stephanie. In 8 years the sum of their ages will be 85. How old is Jared now?
Answer by Boreal(15235)

Stephanie=x
Jared is x+7
In 8 years x+15 + x+8=85
2x+23=85
2x=62
x=31
x+7=38
Jared is now 38 ANSWER
Stephanie is 31
In 8 years, they will be 46 and 39, and that adds to 85.